OpenStudy (anonymous):

can u do me this equation integral of 1/cosx

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}1/cosxdx\]

OpenStudy (dumbcow):

this one can be tricky rewrite it by multiplying by sec+tan on top and bottom \[\rightarrow \frac{1}{\cos x}*\frac{\sec x +\tan x}{\sec x +\tan x} = \frac{\sec^{2} x + \sec x \tan x}{\sec x +\tan x}\] make substitution u = sec x + tan x du = sec *tan + sec^2 dx \[\rightarrow \int\limits_{-1}^{1} \frac{1}{u} du = \ln(u) = \ln(\sec x +\tan x)\]